# Banach algebras associated to twisted étale groupoids: simplicity and pure infiniteness

### v2

@@ Abstract.
We define reduced and essential Banach algebras associated to a twisted étale (not necessarily Hausdorff) groupoid $(\mathcal{G},\mathcal{L})$ and extend some fundamental results from $C^{*}$-algebras to this context. We prove that for topologically free groupoids the associated essential Banach algebras have the ideal interesection property, and thus such an algebra is simple if and only if the groupoid is minimal. We give conditions under which reduced algebras are essential (for example Hausdorffness of $\mathcal{G}$ is sufficient). This in particular solves the simplicity problem posed by Gardella–Lupini for $L^{p}$-operator algebras associated to $\mathcal{G}$. In addition, using either the $n$-filling or locally contracting condition we give pure infiniteness criteria for essential simple Banach algebras associated to $(\mathcal{G},\mathcal{L})$. This extends the corresponding $C^{*}$-algebraic results that were previously known to hold in the untwisted Hausdorff case. The results work nicely, and allow for characterization of the generalized intersection property, in the realm of $L^{P}$-operator algebras where $P\subseteq[1,\infty]$ is a non-empty set of parameters. Such algebras cover in particular $L^{p}$-operator algebras, for $p\in[1,\infty]$, and their Banach *-algebra versions.
@@ 1. Introduction
The characterization of simplicity of the crossed product of a discrete group action on a topological space, [KT90], [AS93], is one of the most fundamental facts from the theory of $C^{*}$-algebras associated to dynamical systems. The dynamical conditions appearing there: topological freeness and minimality, are easy to check, applicable to a number of actions, and most importantly they are also sufficient for simplicity of twisted reduced crossed products. This classical result was generalized to twisted actions of second countable Hausdorff groupoids in [Ren91, Corollary 4.6], though it seemed to gain more attention when phrased in terms of second countable Hausdorff étale groupoids in [Exe11], [BCFS14]. One boost came from Renault’s theory of Cartan inclusions [Ren08], as they are modeled by twisted étale Hausdorff groupoids satisfying a version of topological freeness for groupoids, which in the separable case is topological principality. Since the example of Skandalis from the Appendice to [Ren91], see also [Exe11], it was unclear how to formulate simplicity criteria in the non-Hausdorff case. On the other hand, a number of important constructions lead to non-Hausdorff étale groupoids, cf. [Nek09], [Kat08], [EP16], [EP17], and recently, after a significant breakthrough obtained in [CEPSS19], [KM21], the number of papers on this topic has increased rapidly. In this article we present a cleaned version of this modern theory. The main achievement is that we develop a well-oiled machinery that works not only for $C^{*}$-algebras but in a much broader setting of Banach algebras. By removing a few important stumbling blocks in this setting, such as simplicity problem posed in [GL17], this should open the door to many new applications and research directions connecting the current theories of $C^{*}$-algebras, Banach algebras and harmonic analysis.
@@
Our initial motivation came from Phillips’s program to develop a theory of $L^{p}$-operator algebras, i.e. Banach algebras that can be isometrically represented on a Banach space $L^{p}(\mu)$ for some measure $\mu$, see [Gar21] and/or [Phi12], [Phi13a], [Phi13b], [GT15], [HP15], [GL17], [Phi19], [CR19], [CGT24], [AO22], [HO23], [CMR23] and the references there. In particular, in our previous paper [BKM23] we presented a thorough study of representations of the universal Banach algebras $F^{S}(\mathcal{G},\mathcal{L})$ associated with a twisted étale groupoid $(\mathcal{G},\mathcal{L})$, which are graded by some fixed unital inverse semigroup $S$ of bisections of $\mathcal{G}$. Representations of $F^{S}(\mathcal{G},\mathcal{L})$ on $L^{p}$-spaces do not see this inverse semigroup. In fact for any $p\in[1,\infty]$ the universal $L^{p}$-operator algebras $F^{p}(\mathcal{G},\mathcal{L})$ are graded by the semigroup $\operatorname{Bis}(\mathcal{G})$ of all bisections of $\mathcal{G}$. For $p\neq 2$ their representations can be described in terms of spatial partial isometries [BKM23], and $F^{2}(\mathcal{G},\mathcal{L})$ coincides with the usual universal $C^{*}$-algebra $C^{*}(\mathcal{G},\mathcal{L})$ associated to $(\mathcal{G},\mathcal{L})$ (or a real version of it). To cover other relevant Banach algebras appearing in the literature, such as group Banach algebras [GT15], [LY17], [Phi19], Banach algebra crossed products [DDW11], [BK24], Cuntz Banach algebras [DH22], Roe type Banach algebras [SW17] or *-Banach algebras associated to groupoids [AO22], [HO23], in general we consider Banach algebras $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ that are completions of the underlying convolution algebra $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ of quasi-continuous sections in some seminorm $\|\cdot\|_{\mathcal{R}}$ bounded by the norm coming from some $S$. We view the twist $\mathcal{L}$ over $\mathcal{G}$ as a Fell line bundle over $\mathbb{F}=\mathbb{R},\mathbb{C}$, and the reader not interested in twists may simply ignore it.
@@
In this paper, we turn to study of simplicity and pure infiniteness of algebras of the form $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$. This requires introducing reduced groupoid Banacha algebras of $(\mathcal{G},\mathcal{L})$ and in the non-Hausdorff case also essential groupoid Banach algebras of $(\mathcal{G},\mathcal{L})$. As it is simpler, we discuss first the case when the étale groupoid $\mathcal{G}$ is Hausdorff. This holds if and only if its locally compact Hausdorff unit space $X:=\mathcal{G}^{(0)}$ is a closed subset of $\mathcal{G}$. Then the convolution algebra $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})=C_{c}(\mathcal{G},\mathcal{L})$ consists of compactly supported continuous sections. We say that a Banach algebra $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ is reduced if it is a completion of $C_{c}(\mathcal{G},\mathcal{L})$ such that the inclusion map $C_{c}(\mathcal{G},\mathcal{L})\subseteq C_{0}(\mathcal{G},\mathcal{L})$ extends to an injective contractive map $j_{\mathcal{R}}:F_{\mathcal{R}}(\mathcal{G},\mathcal{L})\to C_{0}(\mathcal{G},\mathcal{L})$ which isometric on $C_{0}(X)$, and so $C_{0}(X)$ is naturally a subalgebra of $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$. This definition covers reduced group algebras [Phi19], reduced crossed products [BK24], as well as the reduced $L^{p}$-algebras $F^{p}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$, $p\in[1,\infty]$ studied in [GL17], [CGT24], [AO22], [HO23], [BKM23]. Following [BK24], we generalize $L^{p}$-operator algebras to $L^{P}$-operator algebras where $P\subseteq[1,\infty]$ is a non-empty set of parameters. In particular, we introduce the full $F^{P}(\mathcal{G},\mathcal{L})$ and reduced $F^{P}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ groupoid $L^{P}$-algebras.
@@
The main difficulty is to establish the relationship between topological freeness of $\mathcal{G}$, as defined in [KM21], and the intersection property for $C_{0}(X)\subseteq F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$, which means that $I\cap C_{0}(X)\neq\{0\}$ for every non-zero (closed two-sided) ideal $I$ in $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$; equivalentely, every representation of $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ which is injective on $C_{0}(X)$ is injective on $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$. We achieve this by extracting the needed pieces and establishing the relationships between the canonical conditional expectation onto $C_{0}(X)$ and the map $j_{\mathcal{R}}$. Here the main idea is to extrapolate the conditional expectation through bisections and the crucial technical tool is the pinching lemma (Lemma 5.8). For a Hausdorff groupoid topological freeness is equivalent to effectivness, i.e. the interior of the isotropy bundle being empty. In particular, combining Proposition 5.11 and Theorems 5.10, 5.16, 5.13 we have the following result:
@@
Let $(\mathcal{G},\mathcal{L})$ be a twisted groupoid where $\mathcal{G}$ is Hausdorff, and denote by $X$ its unit space. Let $P\subseteq[1,\infty]$ be any non-empty set of parameters and let $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ be any reduced twisted groupoid Banach algebra which is graded by some unital inverse semigroup $S$ of bisections covering $\mathcal{G}$. The following conditions are equivalent:
@@
If the above equivalent conditions hold then $C_{0}(X)\subseteq F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ has the intersection property.
The equivalence between (1) and (3) is a far reaching generalization of the Kawamura–Tomiyama theorem [KT90], [AS93], see also [KM21, Theorem 7.29]. Condition (4) is a special case of (3) because for $P\subseteq\{1,\infty\}$ we have $F^{P}(\mathcal{G},\mathcal{L})=F^{P}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ – these algebras are both full and reduced, see [BKM23]. Condition (5) is reminiscent of the recently discovered fact that topological freeness (aperiodicity) is equivalent to the intersection property for all intermediate $C^{*}$-algebras between $C_{0}(X)$ and the reduced $C^{*}$-algebra $C^{*}_{\operatorname{r}}(\mathcal{G},\mathcal{L})=F_{\operatorname{r}}(\mathcal{G},\mathcal{L})$, see [KM22, Theorem 7.3]. As an immediate corollary we get the following simplicity criteria (where by simplicity of a Banach algebra we mean lack of non-trivial closed two-sided ideals):
@@
Corollary B applies to reduced algebras $F^{P}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ for $P\subseteq[1,\infty]$. In particular, it solves in positive the problem posed in [GL17, Problem 8.2], which asked whether minimality and topological principality of a (second countable Hausdorff) groupoid $\mathcal{G}$ implies simplicity of $F^{p}_{\operatorname{r}}(\mathcal{G})$ for $p\in(1,\infty)$. Combining Theorem A with [GL17, Theorem 6.19] we get the following characterization of simplicity of full $L^{P}$-operator algebras for amenable (untwisted) groupoids.
@@
If $\mathcal{G}$ is an amenable second countable locally compact Hausdorff étale groupoid, then for non-empty $P\subseteq[1,\infty]$, then
@@
In the non-Hausdorff setting we generalize the above results by replacing reduced Banach algebras with essential ones. Reduced algebras $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ are completions of $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ equipped with a canonical injective contractive map $j_{\mathcal{R}}:F_{\mathcal{R}}(\mathcal{G},\mathcal{L})\to\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$ where $\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$ is the completion of $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ in $\|\cdot\|_{\infty}$. While by an essential Banach algebra of $(\mathcal{G},\mathcal{L})$ we mean a Hausdorff completion $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ of $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ equipped with an injective contractive map $j_{\mathcal{R}}^{\operatorname{e}}:F_{\mathcal{R}}(\mathcal{G},\mathcal{L})\to\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})/\mathfrak{M}_{0}(\mathcal{G},\mathcal{L})$, where $\mathfrak{M}_{0}(\mathcal{G},\mathcal{L})$ is the subspace of $\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$ consiting of functions with meager open support. Below we phrase these definitions in terms of existence and faithfulness of canonical generalized expectations, which makes clear that they are consistent with existing $C^{*}$-algebraic notions. Every reduced Banach algebra can be quotiened by an ideal to get an essential Banach algebra. There are cases when this ideal is zero, for instance when $\mathcal{G}$ is ample and every compact open set in $\mathcal{G}$ is regular (see Lemma 4.8 below).
Finally we introduce the notion of pure infiniteness for simple (not necessarily unital) Banach algebras, which is consistent with the previous notions of this type (see Definition 6.1 and Proposition 6.3). We modify the classical conditions of locally contracting [A-D97], and $n$-filling [JR00], [Suz17], to include the inverse semigroup $S$ of bisections in $\mathcal{G}$ (Definitions 6.56.8). Combining Theorems 5.10, 6.7 and 6.13 we get:
@@
Groupoid $C^{*}$-algebras are always graded by the full inverse semigroup of bisections $\operatorname{Bis}(\mathcal{G})$. Already in this context the above pure infiniteness criteria are new, as until now they were known to work only in the Hausdorff untwisted case. In general we need to assume local contractiveness or $n$-filling with respect to the inverse semigroup $S(\mathcal{L})$ consisting of bisections where the bundle $\mathcal{L}$ is trivial. But when the bundle $\mathcal{L}$ is trivial, equivalently $\mathcal{L}$ is coming from a groupoid cocycle, we do not need to mention inverse semigroups at all, and we get for instance:
@@
When $\mathcal{G}$ can be covered by a countable family of open bisections then the quotient $F^{P}_{\operatorname{e}}(\mathcal{G},\mathcal{L})$ in the above corollary has a natural injective representation on a direct sum of $L^{p}$-spaces, $p\in P$, and it is natural to view the corresponding Banach algebra as the essential $L^{P}$-groupoid algebra of $(\mathcal{G},\mathcal{L})$ (see Corollary 4.19).
We illustrate our results by using them to generalize and extend the main results of [BK24] from global to partial actions (Subsection 7.1). We present a general groupoid picture for Roe-type Banach algebras of uniformly locally finite coarse structures with coefficients in the compact operators on some Banach space. Such algebras occur naturally, for instance in [SW13], [SW17], [CL18], in connection with criteria for Fredholmness or rigidity. They generalize standard uniform Roe and Roe $C^{*}$-algebras, and the underlying groupoid is always principal so that the equivalent conditions in Theorem A hold (see Subsection 7.2). We introduce reduced and essential twisted tight $L^{P}$-algebras associated to inverse semigroups, and phrase our results in terms of the inverse semigroup. This extends, and sometimes improves, parts of [Exe08], [EP16] where the untwisted $C^{*}$-algebraic case was considered. It also covers all $L^{P}$-groupoid algebras for twisted ample groupoids (Subsection 7.3). Using groupoid methods we extend the main results of [CMR23] from $L^{p}$-graph algebras to $L^{P}$-graph algebras (Subsection 7.4). Finally, we generalize Exel–Pardo algebras [EP17] associated to self-similar actions of groups on graphs to $L^{P}$-operator algebras, and discuss how the recent non-trivial results [EP17], [CEPSS19] extend to this setting (Subsection 7.5). The groupoid modeling these algebras as a rule is non-Hausdorff. Usually the $L^{p}$-algebras for different $p$ are not isomorphic, thus the above results allow us to produce lots of essentially different simple Banach algebras.
The paper is organized as follows. After short preliminaries where we recall some basic facts and notions from [BKM23], in Section 3 we introduce reduced groupoid Banach algebras. In Section 4 we introduce exotic and essential groupoid Banach algebras. Section 5 is devoted to topological freeness and its relationship with other conditions. Section 6 discusses our pure infiniteness results and Section 7.1 presents applications and examples.

### v3

@@ Abstract.
We define reduced and essential Banach algebras associated to a twisted étale (not necessarily Hausdorff) groupoid $(\mathcal{G},\mathcal{L})$ and extend some fundamental results from $C^{*}$-algebras to this context. We prove that for topologically free groupoids the associated essential Banach algebras have the ideal intersection property, and thus such an algebra is simple if and only if the groupoid is minimal. We give conditions under which reduced algebras are essential (for example Hausdorffness of $\mathcal{G}$ is sufficient). This in particular solves the simplicity problem posed by Gardella–Lupini for $L^{p}$-operator algebras associated to $\mathcal{G}$. In addition, using either the $n$-filling or locally contracting condition we give pure infiniteness criteria for essential simple Banach algebras associated to $(\mathcal{G},\mathcal{L})$. This extends the corresponding $C^{*}$-algebraic results that were previously known to hold in the untwisted Hausdorff case. The results work nicely, and allow for characterization of the generalized intersection property, in the realm of $L^{P}$-operator algebras where $P\subseteq[1,\infty]$ is a non-empty set of parameters. Such algebras cover in particular $L^{p}$-operator algebras, for $p\in[1,\infty]$, and their Banach *-algebra versions.
@@ 1. Introduction
The characterization of simplicity of the crossed product of a discrete group action on a topological space, [KT90], [AS93], is one of the most fundamental facts from the theory of $C^{*}$-algebras associated to dynamical systems. The dynamical conditions appearing there, topological freeness and minimality, are easy to check, applicable to a number of actions, and, most importantly, they are also sufficient for simplicity of twisted reduced crossed products. This classical result was generalized to twisted actions of second countable Hausdorff groupoids in [Ren91, Corollary 4.6], though it seemed to gain more attention when phrased in terms of second countable Hausdorff étale groupoids in [Exe11], [BCFS14]. One boost came from Renault’s theory of Cartan inclusions [Ren08], as they are modeled by twisted étale Hausdorff groupoids satisfying a version of topological freeness for groupoids, which in the separable case is topological principality. Since the example of Skandalis from the Appendice to [Ren91], see also [Exe11], it was unclear how to formulate simplicity criteria in the non-Hausdorff case. On the other hand, a number of important constructions lead to non-Hausdorff étale groupoids, cf. [Nek09], [Kat08], [EP16], [EP17], and recently, after a significant breakthrough obtained in [CEPSS19], [KM21], the number of papers on this topic has increased rapidly. In this article, we present a cleaned version of this modern theory. The main achievement is that we develop a well-oiled machinery that works not only for $C^{*}$-algebras but in the much broader setting of Banach algebras. By removing a few important stumbling blocks in this setting, such as the simplicity problem posed in [GL17], this should open the door to many new applications and research directions connecting the current theories of $C^{*}$-algebras, Banach algebras and harmonic analysis.
@@
Our initial motivation came from Phillips’s program to develop a theory of $L^{p}$-operator algebras, that is, Banach algebras that can be isometrically represented on a Banach space $L^{p}(\mu)$ for some measure $\mu$, see [Gar21] and/or [Phi12], [Phi13a], [Phi13b], [GT15], [HP15], [GL17], [Phi19], [CR19], [CGT24], [AO22], [HO23], [CMR23] and the references there. In particular, in our previous paper [BKM25] we presented a thorough study of representations of the universal Banach algebras $F^{S}(\mathcal{G},\mathcal{L})$ associated with a twisted étale groupoid $(\mathcal{G},\mathcal{L})$, which are graded by some fixed unital inverse semigroup $S$ of bisections of $\mathcal{G}$. Representations of $F^{S}(\mathcal{G},\mathcal{L})$ on $L^{p}$-spaces do not see this inverse semigroup. In fact for any $p\in[1,\infty]$ the universal $L^{p}$-operator algebras $F^{p}(\mathcal{G},\mathcal{L})$ are graded by the semigroup $\operatorname{Bis}(\mathcal{G})$ of all bisections of $\mathcal{G}$. For $p\neq 2$, their representations can be described in terms of spatial partial isometries [BKM25], and $F^{2}(\mathcal{G},\mathcal{L})$ coincides with the usual universal $C^{*}$-algebra $C^{*}(\mathcal{G},\mathcal{L})$ associated to $(\mathcal{G},\mathcal{L})$ (or a real version of it). To cover other relevant Banach algebras appearing in the literature, such as group Banach algebras [GT15], [LY17], [Phi19], Banach algebra crossed products [DDW11], [BK24], Cuntz Banach algebras [DH22], Roe type Banach algebras [SW17] or *-Banach algebras associated to groupoids [AO22], [HO23], in general we consider Banach algebras $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ that are completions of the underlying convolution algebra $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ of quasi-continuous sections in some seminorm $\|\cdot\|_{\mathcal{R}}$ bounded by the norm coming from some $S$. We view the twist $\mathcal{L}$ over $\mathcal{G}$ as a Fell line bundle over $\mathbb{F}=\mathbb{R},\mathbb{C}$, and the reader not interested in twists may simply ignore it.
@@
In this paper, we turn to the study of simplicity and pure infiniteness of algebras of the form $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$. This requires introducing reduced groupoid Banach algebras of $(\mathcal{G},\mathcal{L})$ and in the non-Hausdorff case also essential groupoid Banach algebras of $(\mathcal{G},\mathcal{L})$. As it is simpler, we discuss first the case when the étale groupoid $\mathcal{G}$ is Hausdorff. This holds if and only if its locally compact Hausdorff unit space $X:=\mathcal{G}^{(0)}$ is a closed subset of $\mathcal{G}$. Then the convolution algebra $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})=C_{c}(\mathcal{G},\mathcal{L})$ consists of compactly supported continuous sections. We say that a Banach algebra $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ is reduced if it is a completion of $C_{c}(\mathcal{G},\mathcal{L})$ such that the inclusion map $C_{c}(\mathcal{G},\mathcal{L})\subseteq C_{0}(\mathcal{G},\mathcal{L})$ extends to an injective contractive map $j_{\mathcal{R}}:F_{\mathcal{R}}(\mathcal{G},\mathcal{L})\to C_{0}(\mathcal{G},\mathcal{L})$ which isometric on $C_{0}(X)$, and so $C_{0}(X)$ is naturally a subalgebra of $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$. This definition covers reduced group algebras [Phi19], reduced crossed products [BK24], as well as the reduced $L^{p}$-algebras $F^{p}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$, $p\in[1,\infty]$ studied in [GL17], [CGT24], [AO22], [HO23], [BKM25]. Following [BK24], we generalize $L^{p}$-operator algebras to $L^{P}$-operator algebras where $P\subseteq[1,\infty]$ is a non-empty set of parameters. In particular, we introduce the full $F^{P}(\mathcal{G},\mathcal{L})$ and reduced $F^{P}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ groupoid $L^{P}$-algebras.
@@
The main difficulty is to establish the relationship between topological freeness of $\mathcal{G}$, as defined in [KM21], and the intersection property for $C_{0}(X)\subseteq F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$, which means that $I\cap C_{0}(X)\neq\{0\}$ for every non-zero (closed two-sided) ideal $I$ in $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$; equivalentely, every representation of $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ which is injective on $C_{0}(X)$ is injective on $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$. We achieve this by establishing the relationships between the canonical conditional expectation onto $C_{0}(X)$ and the map $j_{\mathcal{R}}$. Here the main idea is to extrapolate the conditional expectation through bisections and the crucial technical tool is the pinching lemma (Lemma 5.8). For a Hausdorff groupoid, topological freeness is equivalent to effectiveness, that is, the interior of the isotropy bundle being only the unit space. In particular, combining Proposition 5.11 and Theorems 5.10, 5.16, 5.13 we have the following result:
@@
Let $(\mathcal{G},\mathcal{L})$ be a twisted groupoid where $\mathcal{G}$ is a locally compact Hausdorff étale groupoid with unit space $X$. Let $P\subseteq[1,\infty]$ be any non-empty set of parameters and let $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ be any reduced twisted groupoid Banach algebra, which is graded by some unital inverse semigroup $S$ of bisections covering $\mathcal{G}$. The following conditions are equivalent:
@@
If the above equivalent conditions hold, then $C_{0}(X)\subseteq F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ has the intersection property.
The equivalence between (1) and (3) is a far reaching generalization of the Kawamura–Tomiyama theorem [KT90], [AS93], see also [KM21, Theorem 7.29]. Condition (4) is a special case of (3) because for $P\subseteq\{1,\infty\}$ we have $F^{P}(\mathcal{G},\mathcal{L})=F^{P}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ – these algebras are both full and reduced, see [BKM25]. Condition (5) is reminiscent of the recently discovered fact that topological freeness (aperiodicity) is equivalent to the intersection property for all intermediate $C^{*}$-algebras between $C_{0}(X)$ and the reduced $C^{*}$-algebra $C^{*}_{\operatorname{r}}(\mathcal{G},\mathcal{L})=F_{\operatorname{r}}(\mathcal{G},\mathcal{L})$, see [KM22, Theorem 7.3]. As an immediate corollary we get the following simplicity criteria (where by simplicity of a Banach algebra we mean lack of non-trivial closed two-sided ideals):
@@
Corollary B applies to reduced algebras $F^{P}_{\operatorname{r}}(\mathcal{G},\mathcal{L})$ for $P\subseteq[1,\infty]$. In particular, it solves in positive the problem posed in [GL17, Problem 8.2], which asked whether minimality and topological principality of a (second countable Hausdorff) groupoid $\mathcal{G}$ implies simplicity of $F^{p}_{\operatorname{r}}(\mathcal{G})$ for $p\in(1,\infty)$. Combining Theorem A with [GL17, Theorem 6.19], we get the following characterization of simplicity of full $L^{P}$-operator algebras for amenable (untwisted) groupoids.
@@
If $\mathcal{G}$ is an amenable second countable locally compact Hausdorff étale groupoid, then for any non-empty set $P\subseteq[1,\infty]$ we have that
@@
In the non-Hausdorff setting, we generalize the above results by replacing reduced Banach algebras with essential ones. Reduced algebras $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ are completions of $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ equipped with a canonical injective contractive map $j_{\mathcal{R}}:F_{\mathcal{R}}(\mathcal{G},\mathcal{L})\to\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$ where $\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$ is the completion of $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ in $\|\cdot\|_{\infty}$ (see Definition 3.1). By an essential Banach algebra of $(\mathcal{G},\mathcal{L})$ we mean a Hausdorff completion $F_{\mathcal{R}}(\mathcal{G},\mathcal{L})$ of $\mathfrak{C}_{c}(\mathcal{G},\mathcal{L})$ equipped with an injective contractive map $j_{\mathcal{R}}^{\operatorname{e}}:F_{\mathcal{R}}(\mathcal{G},\mathcal{L})\to\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})/\mathfrak{M}_{0}(\mathcal{G},\mathcal{L})$, where $\mathfrak{M}_{0}(\mathcal{G},\mathcal{L})$ is the subspace of $\mathfrak{C}_{0}(\mathcal{G},\mathcal{L})$ consisting of functions with meager open support (see Definition 4.9). Below we phrase these definitions in terms of existence and faithfulness of canonical generalized expectations, which makes clear that they are consistent with existing $C^{*}$-algebraic notions. Every reduced Banach algebra can be quotiented by an ideal to get an essential Banach algebra. There are cases when this ideal is zero, for instance when $\mathcal{G}$ is ample and every compact open set in $\mathcal{G}$ is regular (see Lemma 4.8 below).
Finally, we introduce the notion of pure infiniteness for simple (not necessarily unital) Banach algebras, which is consistent with the previous notions of this type (see Definition 6.1 and Proposition 6.3). We modify the classical conditions of locally contracting [A-D97], and $n$-filling [JR00], [Suz17], to include the inverse semigroup $S$ of bisections in $\mathcal{G}$ (Definitions 6.56.8). Combining Theorems 5.10, 6.7 and 6.13 we get:
@@
Groupoid $C^{*}$-algebras are always graded by the full inverse semigroup of bisections $\operatorname{Bis}(\mathcal{G})$. Already in this context, the above pure infiniteness criteria are new, as until now they were known to work only in the Hausdorff untwisted case. In general, we need to assume local contractiveness or $n$-filling with respect to the inverse semigroup $S(\mathcal{L})$ consisting of bisections where the bundle $\mathcal{L}$ is trivial. But when the bundle $\mathcal{L}$ is trivial, equivalently $\mathcal{L}$ is coming from a groupoid cocycle, we do not need to mention inverse semigroups at all, and we get for instance:
@@
If $\mathcal{G}$ can be covered by a countable family of open bisections, then the quotient $F^{P}_{\operatorname{e}}(\mathcal{G},\mathcal{L})$ in the above corollary has a natural injective representation on a direct sum of $L^{p}$-spaces, $p\in P$, and it is natural to view the corresponding Banach algebra as the essential $L^{P}$-groupoid algebra of $(\mathcal{G},\mathcal{L})$ (see Corollary 4.19).
We illustrate our results by using them to generalize and extend the main results of [BK24] from global to partial actions (Subsection 7.1). We present a general groupoid picture for Roe-type Banach algebras of uniformly locally finite coarse structures with coefficients in the compact operators on some Banach space. Such algebras occur naturally, for instance in [SW13], [SW17], [CL18], in connection with criteria for Fredholmness or rigidity. They generalize standard uniform Roe and Roe $C^{*}$-algebras, and the underlying groupoid is always principal so that the equivalent conditions in Theorem A hold (see Subsection 7.2). We introduce reduced and essential twisted tight $L^{P}$-algebras associated to inverse semigroups, and phrase our results in terms of the inverse semigroup. This extends, and sometimes improves, parts of [Exe08], [EP16] where the untwisted $C^{*}$-algebraic case was considered. It also covers all $L^{P}$-groupoid algebras for twisted ample groupoids (Subsection 7.3). Using groupoid methods we extend the main results of [CMR23] from $L^{p}$-graph algebras to $L^{P}$-graph algebras (Subsection 7.4). Finally, we generalize Exel–Pardo algebras [EP17] associated to self-similar actions of groups on graphs to $L^{P}$-operator algebras, and discuss how the recent non-trivial results [EP17], [CEPSS19] extend to this setting (Subsection 7.5). The groupoid modeling these algebras, as a rule, is non-Hausdorff. Usually the $L^{p}$-algebras for different $p$ are not isomorphic, thus the above results allow us to produce lots of essentially different simple Banach algebras.
The paper is organized as follows. After short preliminaries where we recall some basic facts and notions from [BKM25], in Section 3 we introduce reduced groupoid Banach algebras. In Section 4 we introduce exotic and essential groupoid Banach algebras. Section 5 is devoted to topological freeness and its relationship with other conditions. Section 6 discusses our pure infiniteness results and Section 7 presents applications and examples.